OpenStudy (anonymous):

If the parent function f(x) = (2x – 3)³ is transformed to g(x) = (-2x + 3)³, which type of transformation occurs?

geerky42 (geerky42):

Looks like f(x) was multiplied by -1 to get to g(x).

geerky42 (geerky42):

\((2x-3)^3 \quad\Longrightarrow\quad-(2x-3)^3\\~\\\phantom{(2x-3)^3 \quad\Longrightarrow~}=(-1)^3(2x-3)^3\\~\\\phantom{(2x-3)^3 \quad\Longrightarrow~}=[(-1)(2x-3)]^3\\~\\\phantom{(2x-3)^3 \quad\Longrightarrow~}=(-2x+3)^3\)

OpenStudy (anonymous):

i just need to know vertical shift/refection or horizontal shift/reflection

OpenStudy (welshfella):

this is a reflection in which axis do you think?

OpenStudy (anonymous):

I think I would graph it.